OpenStudy (anonymous):

George and Carmen went on a bicycle trip. The took a bus to their starting point, and then biked the rest. They traveled 350 kilometers in total, and they biked 80 kilometers more than they were bussed Let x=kilometers traveled by bike and y=kilometers traveled by bus Find how many kilometers they traveled by bike?

OpenStudy (anonymous):

hey um i am open!

OpenStudy (shamim):

x+y=350 km Right?

OpenStudy (anonymous):

yeah

OpenStudy (shamim):

x=y+80 Right?

OpenStudy (anonymous):

yeah

OpenStudy (shamim):

Frm my 2 equations tell me x=? y=?

OpenStudy (anonymous):

thanks sooo much

OpenStudy (shamim):

Just put the value of x frm my second equation into first equation

OpenStudy (shamim):

I think u r not able to find out x=? y=?

OpenStudy (shamim):

Ok c x+y=350 But x=y+80 So y+80+y=350 y=?

OpenStudy (shamim):

Plz try it a little !

OpenStudy (shamim):

Or simply tell me u r not getting y=?

OpenStudy (anonymous):

I am still not getting y please help

OpenStudy (anonymous):

I Got It It WAS 350-80=x X/2=y y+80=the answer